I am trying to come up with an equation for a parabola and I cannot seem to figure it out. I have an algebra 2 project where I have to write a paper on a problem that relates to my life. I want to find out the parabola of a corner kick in soccer. The corner is 25 yards away from the player the ball is being kicked to. The player is 5 foot 6 inches tall. I want the ball to be kicked from the corner to the players head. How could I write an equation for that?
Answer by venugopalramana(3286)

I am trying to come up with an equation for a parabola and I cannot seem to figure it out. I have an algebra 2 project where I have to write a paper on a problem that relates to my life. I want to find out the parabola of a corner kick in soccer. The corner is 25 yards away from the player the ball is being kicked to. The player is 5 foot 6 inches tall. I want the ball to be kicked from the corner to the players head. How could I write an equation for that?
STD.EQN.OF PARABOLA TAKING YOUR POSITION AS VERTEX (0,0)AND THE LINE BETWEEN YOU AND THE RECEIVER AS AXIS IS
Y^2=4AX
THE RECEIVER'S POSITION IS (25*3=75',5.5')IS WHERE THE PARABOLA SHOULD PASS THROUGH...SO
5.5^2=4A*75=300A
A=5.5*5.5/300=0.10083333
SO EQN. OF PARABOLA IS Y^2=0.403333X